# An Improved Two-Step Attack on CRYSTALS-Kyber

### v1

@@ 3.2. Modified Leakage Model for CPA
Modeling the power leakage is the basis for a CPA attack, and its accuracy directly determines the success rate of the attack. HW model counts the number of “1” in intermediate values, which is one of the most representative linear power models. Similarly, HD model records the amount of ”1$\rightarrow$”0” or ”0$\rightarrow$”1” transitions as power leakage. The difference of HW and HD model is given as:
(6) $HD\left(r_{0},s_{1}\right)=HW\left(r_{0}\oplus s_{1}\right).$
Typically, when simulating the energy consumption of a microprocessor, the HW model is commonly employed (zhao2023side, ), as is the case in this particular study.
The arithmetic logic units (ALUs) and multiplier units (MU) in ARM processors are often optimized for two’s complement operations. Most arithmetic and logical instructions in the processor’s instruction set are designed based on two’s complement calculations. Therefore, we fine-tune the HW power leakage model according to this property to directly avoid the complementary false positive. We compute the HW of the complement of the intermediate instead of directly computing the HW of the intermediate, as shown in Listing 2, where the function calculate_complement computes the $16bit$ complement. The modified power leakage model can map to the hypothetical power consumption values more accurately, thus reducing the number of power traces required for CPA attacks. As show in Fig. 2, correct secret key coefficient in the original HW leakage model is drowned in all the guessed key coefficients, while the coefficient can be picked out in the modified model effectively by using a small number of 15 power traces.
Figure 2. When $D=15$ during a $basemul$ call, PCC values for all guessed keys of the original HW model (top) and the modified HW model (bottom).
@@ 3.3. Optional Kendall’s tau for CPA
The goal of this improvement is to find the correct key coefficients using 15 or even fewer power traces while ensuring the efficiency and success rate of the attack. In scenarios where the number of power traces is limited, the calculated PCC value across all guessed keys tends to be exceptionally high. Consequently, it becomes challenging to accurately recover even a portion of the secret key coefficients using solely the traditional CPA attack, and attempting to reconstruct 128 or the entire set of secret key coefficients becomes even more daunting.
Our solution is as follows. After computing the PCC correlation matrix for each $basemul$, we set the threshold to 0.9 or even higher, adjusting it based on the values of $\mathbf{P}$. For each $basemul$, if there are more than $f$ correlations exceeding the threshold, we consider them as candidate coefficients $\mathbf{k_{h}}=(k_{h_{0}},...,k_{h_{f-1})}$. We then map the hypothetical intermediate values of the $f$ candidate coefficients to candidate hypothetical power consumption value matrix $\mathbf{H^{\prime}}_{D\times f}$. Next, we calculate the Kendall’s tau between the hypothetical power value matrix $\mathbf{H^{\prime}}$ and the actual power matrix $\mathbf{T}$. For each column of $\mathbf{H^{\prime}}$, we compute the Kendall’s tau with each column of the actual power matrix. Finally, we obtain a $f\times T$ Kendall coefficient matrix $\mathbf{K_{d}}$, from which we select the candidate key with the highest correlation coefficient as the correct key coefficient for that $basemul$. On the other hand, we skip the cases that have higher PCC correlations and smaller candidates.
PCC is used in the traditional CPA attack to calculate only the linear correlation between the actual power consumption and the hypothetical power consumption, which is not sufficient. With only a few power traces, it is possible to directly select some of the correct key coefficients even if only Kendall’s tau is calculated, but the calculation time of Kendall’s tau is 12×12$\times$12 × longer than that of PCC. Therefore we combine Kendall’s tau with PCC. After calculating PCC, the guessed keys coefficient exceeding a certain threshold is selected as the coefficients of candidate keys. Then, the Kendall’s tau values of these coefficients of candidate keys are calculated so that the coefficient with the higher value is accepted. In this way, false positives can be quickly eliminated.
3.4. Trail-and-Error Lattice Attack after CPA
As mentioned earlier, each group of the secret key consists of 128 secret key coefficients including 128 even-index coefficients or 128 odd-index coefficients. As introduced in (kuo2023lattice, ), suppose that an attacker wants to exploit the side-channel leakage point $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$ and has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered.
@@ 3.4. Trail-and-Error Lattice Attack after CPA
Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients. INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients. Similarly, $\mathbf{N}_{B}=[\mathbf{n}_{b_{0}},\ldots,\mathbf{n}_{{b}_{127-sr}}]$, and $\hat{\mathbf{s}}_{iB}=[\hat{\mathbf{s}}_{b_{0}},\ldots,\hat{\mathbf{s}}_{{b}_{127-sr}}]^{\top}$.
@@
In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known. Let $\mathbf{t}=\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$, $\mathbf{A}=-\mathbf{N}_{B}$, and $\mathbf{s}^{\prime}=\hat{\mathbf{s}}_{iB}$. Then, we obtain $\mathbf{t}=\mathbf{A}\cdot\mathbf{s}^{\prime}+\mathbf{s}_{i}\mod q$. This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller. We firstly view it as a bounded distance decoding (BDD)/unique shortest vector problem (uSVP) lattice problem, and specific algorithm to solve the updated LWE problem is given by Algorithm 2 (kannan1987minkowski, ), where the matrixs $\mathbf{BDD}$ and $\mathbf{B}_{kan}$ are summarized as follows:
(7) $\mathbf{BDD}=\left[\begin{array}[]{cc}\mathbf{I}_{sr}&\mathbf{A}^{\prime}\\ \mathbf{0}&q\mathbf{I}_{n-sr}\\ \end{array}\right],$
@@
$\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose. This process yields the shortest vector containing the error vector $\mathbf{s}_{i}^{\top}$, denoted as $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$. Finally, We obtain the actual secret key with a length of 128 as shown in Step 3 of Algorithm 2.
But it is worth noting that the norm of vector $\mathbf{w}$ is $\sqrt{\|s_{i}\|^{2}+1}\approx\sqrt{n}\sigma_{s}$, which must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the uSVP problem to be solved. For Kyber512, when $sr\geq 39$, it is possible to correctly recover the length of 128 odd/even index coefficients through lattice attack. For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered. In other words, the lattice attack can tolerate at most $89/90$ recovered incorrect secret key coefficients for Kyber512 and Kyber768/1024 respectively, where $38/39$ of the coefficient are randomly selected from $128$ recovery coefficients.
We elaborate on the theoretical principles of the lattice attack, which provides a fault-tolerant for recovering the full set of 128 keys. Leveraging this method, we propose a more flexible and versatile approach for its application. The proposed trail-and-error lattice attack is shown in Listing 3, where $n$ represents the number of CPA attacks, and $\mathbb{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trails, respectively.
Listing 3: Trail-and-Error Lattice Attack.
@@
Algorithm 2 Kannan’s embedding technique.
1: Input: An LWE instance $\mathbf{t}=\mathbf{A}\mathbf{s}+\mathbf{s}_{i}\bmod q$
2: Output: the short error vector $\mathbf{s}_{i}\in\mathcal{R}_{q}^{128}$
3:Construct the lattice $\Lambda(\text{BDD})$ generated by matrix $\mathbf{BDD}$;
4:Reduce BDD to the SVP problem by rescaling $\mathbf{BDD}$ to the basis matrix $\mathbf{B}_{kan}$;
5:Use lattice algorithm (LLL or BKZ) to derive the short vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ from $\mathbf{B}_{kan}$;
6:return $\mathbf{s}_{i}$
One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA attacks instead of $128$ CPA attacks. Certainly, the more CPA attacks are carried out, the higher the success rate of this method will be, but its efficiency will also slow down. The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trail-and-error method. It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trail and lead to the overall recovery inefficient.
For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively. Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
Figure 3. The experimental equipment.
@@ 4. Experiments
4.1. Experimental Setup
Our attacks are equally successful for different security levels of Kyber because they use the same $basemul$, whose algorithm is provided by Listing 1. For simplicity, we only take Kyber512 into consideration in the following. We run the reference implementation of Kyber512 from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz. We capture the power traces using a Pico 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe. The sampling rate is set to $500$ MSa/s. The experimental equipment is shown in Fig. 3.
According to the introduction in Section 2.4 that the selection of the intermediate value of CPA attack should be related to the secret key, combined with the analysis mentioned in Section 3.1, we choose $fqmul(k,u_{1})$ as the intermediate value of our attack, and all the intermediate values in the following experiments are the same values.
Figure 4. When $D=15$, the PCC results during a certain $basemul$ call.
4.2. Experiment Results
Step1: CPA Attack
As aforementioned, the CPA attack is divided into two stages: the capturing stage and the modeling computation stage. Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
For the capturing stage, as introduced in Section 3.1, we carry out the CPA attack at the NTT operation which calls $basemul$ functions 128 times. In our CPA experiments, we input random ciphertexts and capture tremendous electromagnetic radiations during those calls.
To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments. The first one is shown in Fig. 2 of Section 3.2 and we can see that the new model can reinforce the PCC value of the correct key. The second one shown in the following is to demonstrate its superiority in identifying the complementary false positive.
@@ 4.2. Experiment Results
Figure 5. When $D=15$, the PCC results (a) and Ken results (b) in a certain $basemul$. When $D=11$, the PCC results (c) and Ken results (d) in the same $basemul$.
@@
When the number of random ciphertexts $D$ is set to 15, the relationship of the maximum absolute values $\left|PCC\right|$ to the guess keys is shown in the above of Fig. 4. We can directly pick out the correct key 2280 and its complementary false positive 1049 (equal to $3329-2280$). This is a general phenomenon. As mentioned in (kuo2023lattice, ), the attacker would run into some problems because the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ both would have high $PCC$ values since $HW(s_{0})$ and $HW(q-s_{0})$ are highly correlated. So, we have modified the leakage model as described in Section 3.2 to escape such false positive. The relationship of the PCC values to the trace points is shown in the bottom of of Fig. 4, where there are 3329 curves and the yellow and red curves present whose guess keys are 2280 and 1049, respectively. It can be seen that the maximum absolute value of PCC of the complementary false positive has a negative sign. We can only focus on the positive PCCs to avoid such false positive.
We briefly explain the reason of this phenomenon in the following. Let $s_{0}$ be the correct one. Then $s_{0}^{\prime}=3329-s_{0}$ represents the false positive. The $fqmul$ function in $basemul$ is an operation that multiplies two numbers and then is reduced by $3329$ in Kyber. The intermediate values for the two guessed key coefficients are summarized as follows:
@@
Hypothetical power consumption values $MHW(r_{1})$ and $MHW(r_{1}^{\prime})$ exhibit a negative correlation, while correct $MHW(r_{1})$ will naturally exhibit a positive correlation. Obviously, we can figure them out by taking their signs into consideration.
@@
To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process. Figure 5 (a) shows the points of the maximum absolute PCC results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates $f$ after PCC is equal to 4. It can be seen that the correlations of four coefficients $564$, 1687, 2010, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to 0.930 and their correlation values are very close. Very close correlations can lead to an inability to distinguish the correct coefficient. This is an another kind of false positives. If the attacker just picks one of these four at random, the accuracy is only about 25%, which is still regarded as a failure. Therefore, we use the method mentioned in Section 3.3 to compensate for such case. First of all, we compute the Kendall coefficient matrix of these four candidate coefficients. It should be noted that the time for this computation is negligible when comparing to the complete PCC computing. Figure 5 (b) shows the points of the four maximum absolute values $\left|Kendall’s\,tau\right|$ to the numbers of power traces. It shows that Kendall’s tau expands the differences between correct coefficient and false positives. As the number of power traces increases, the value of correct coefficient tends to be more stable while those of false positives decrease quickly. We can easily distinguish the correct key 564 when the number of power traces equals 15 and output it as the final result. In another word, an attacker can take the correct coefficient 564 for this $basemul$ call using only 15 power traces. Similarly, we reduce the number of power traces to 11 step by step and implement the PCC and Kendall’s tau computations. The corresponding results are shown in Fig. 5 (c) and (d). It can be seen that the correct key 564 can almost not be recognized. We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections. Note that if we continue reducing the number of power traces, we cannot figure it out any more.
Step2: Trail-and-Error Lattice Attack
We conduct the trail-and-error lattice attack on a sixteen-thread server. The parameter of the $block$ size in the BKZ reduction is set to 50 and the value of $max\_loops$ to 8. We collect the CPAs of the first sixty $basemuls$ calls.
@@
Figure 6 shows the relationships of the success rate and time to the number of power traces representing in the blue curve and groups of histograms, respectively. The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times. It can be seen that when the number of power traces is no less than 15, all the experimental attacks are successful. Those cases with traces number smaller than 15 can be compensated to some degree by running more trails of the lattice attack. Meanwhile, according to the decomposed time histograms in Fig. 6, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA attack is almost unchanged although the traces number is doubled. The main reason could be the different utilization ratios of the multi-threading server. Note that the CPA computations are independently executed with the different power traces while the trail computations of the lattice attack are executed in serial.
Therefore, we can summarize that when the success rate is large enough (close to 1), we can only cost 15 power traces for a guess/call and about 9 minutes to recover the full key on Kyber512. So, the total number of required power traces is $15\times 60=900$, much smaller than that of the state-of-the-art random ciphertext CPA (yang2023chosen, ) which requires $20\times 128=2560$ power traces. If we slightly increasing the traces number of the guess, the required time can be reduced close to 5 minutes, much faster than the recorded 20 minutes reported in (kuo2023lattice, ).

### v2

@@ 3.1. Modified Leakage Model for CPA
The Hamming weight (HW) model counts the number of ”1” in intermediate values and is a common linear power model. ARM processors’ ALUs are optimized for two’s complement operations, so we adjust the HW (called modified Hamming weight, MHW) to account for this by using the HW of the 16-bit complement of the intermediate values. This modification improves the accuracy of the HW model and reduces the number of power traces (denoted as $D$) needed for CPA. As shown inside the blue dotted box in Fig. 2 (a) and (b), the correct key coefficient is obscured using the original model (a) but can be effectively identified with just 15 power traces using the modified model (b).
@@ 3.2. Optional Kendall’s tau for CPA
When the number of power traces is limited, traditional CPA may face challenges due to high PCC values across guessed keys. To address this, we set a PCC threshold (e.g., 0.9) and extract candidate key coefficients with correlations exceeding this threshold. We then map these candidates to a hypothetical power consumption matrix and compute Kendall’s tau between this matrix and the actual power traces. The candidate with the highest Kendall’s tau value is selected as the correct key coefficient, which helps to reduce false positives. This approach combines PCC for initial filtering with Kendall’s tau for more precise selection, despite Kendall’s tau being more computationally demanding.
Figure 2. (a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively. (c) shows the ||||PCC|||| values for all guessed keys, while the blue curve in (d) illustrates the occurrence of false positives.
3.3. Trial-and-Error Lattice Attack after CPA
@@
As introduced in (kuo2023lattice, ), suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered. Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients. INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients. Similarly, $\mathbf{N}_{B}$ and $\hat{\mathbf{s}}_{iB}$ are obtained in the same manner.
@@
In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known. Let $\mathbf{t}=\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$, $\mathbf{A}=-\mathbf{N}_{B}$, and $\mathbf{s}^{\prime}=\hat{\mathbf{s}}_{iB}$. Then, we obtain $\mathbf{t}=\mathbf{A}\cdot\mathbf{s}^{\prime}+\mathbf{s}_{i}\mod q$. This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller. We use Kannan’s embedding technique to solve this problem (kannan1987minkowski, ), where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
@@
$\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose. A vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ of length $129$ is obtained by reduction algorithm, which contains the secret key $\mathbf{s}_{i}^{\top}$ of length 128.
But it is worth noting that the norm of vector $\mathbf{w}$ must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the unique shortest vector problem to be solved. For Kyber512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients. For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
We elaborate on the theoretical foundation of lattice attacks, which can tolerate error coefficients of recovered key and still successfully retrieve all 128 key. Leveraging this method, we propose a more flexible and versatile approach for its application. The proposed trial-and-error lattice attack is shown in Listing 2, where $n$ represents the number of CPA, and $\mathbf{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trials, respectively.
One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA instead of $128$ CPA. Certainly, the more CPA are carried out, the higher the success rate of this method will be, but its efficiency will also slow down. The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trial-and-error method. It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trial and lead to the overall recovery inefficient. For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively. Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
Listing 2: Trail-and-Error Lattice Attack.
@@ 4. Experiments
The proposed attack is applied to Kyber512, and it is noteworthy that our method is also applicable to other security levels due to their use of the same $basemul$. We run the reference implementation from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz. Power traces were captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a sampling rate set to $500$ MSa/s. Based on the analysis in Section 3, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
@@
Figure 3. (a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$. Similarly, (c) and (d) show the PCC and Ken results for the same $basemul$ call when $D=11$.
Step1: CPA
As aforementioned, the CPA is divided into two stages: the capturing stage and the modeling computation stage. Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions for 128 times. In our CPA experiments, we capture random ciphertexts and tremendous data of electromagnetic radiations during those calls.
To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments. The former is shown in Fig. 2 (a) and (b) of Section 3.1 and the modified model can reinforce the PCC value of the correct key. The latter shown in Fig. 2 (c) and (d) is to demonstrate its superiority in identifying the complementary false positive.
@@
When the number of random ciphertexts $D$ is set to $15$, the relationship of the maximum ||||PCC|||| to the guess keys is shown in Fig. 2 (c). We can directly pick out the correct key $2280$ and its complementary false positive $1049$ (equal to $3329-2280$), which is a general phenomenon. In addition, as the number of power traces decreases, the attacker would run into some problems because both the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ would have high and close PCC values, as the HW$(s_{0})$ and HW$(q-s_{0})$ are highly correlated. To escape such false positive, we have modified the leakage model as described in Section 3.1. The correspondence between the PCC values of the correct key coefficient and the false positive one is shown in Fig. 2 (d). It is observed that the positive peak value corresponds to the correct key coefficient and the negative peak value of PCC corresponds to the false positive one. Consequently, to avoid the false positive one is selected, we should focus solely on the positive peak value of PCC. We briefly explain this phenomenon as follows: Let $s_{0}$ be the correct one, and $s_{0}^{\prime}=3329-s_{0}$ be the false positive one. The intermediate values for the two guessed key coefficients are summarized below:
@@
MHW$(r_{1})$ and MHW$(r_{1}^{\prime})$ show a negative correlation, where the correct MHW$(r_{1})$ is expected to be a positive PCC. Thus, considering their signs allows us to distinguish them.
@@
To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process. Fig. 3 (a) shows the points of the maximum ||||PCC|||| results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates after PCC is equal to 4. It can be seen that the correlations of four coefficients $564$, $1687$, $2010$, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to $0.930$ and their correlation values are very close. The phenomenon can lead to an inability to distinguish the correct coefficient. This is another kind of false positives. If the attacker just picks one of the four candidates at random, the accuracy is only about $25$%, which is still regarded as a failure. Therefore, we use the method mentioned in Section 3.2 to compensate for such case. First of all, we compute the Kendall coefficient matrix of these four candidate coefficients. It should be noted that the time of this computation is negligible when compared to the complete PCC computing. Fig. 3 (b) shows the values of the four maximum ||||Kendall’s tau|||| when $D=15$. It shows that Kendall’s tau expands the differences between correct coefficient and false positives. As the number of power traces increases, the value of correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly. We can easily distinguish the correct key $564$ when $D=15$ and output it as the final result. In another word, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces. Similarly, we reduce the number of power traces to $11$ step by step and implement the PCC and Kendall’s tau computations. The corresponding results are shown in Fig. 3 (c) and (d). It can be seen that the correct key $564$ can almost not be recognized. We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections. Note that if we continue reducing the number of power traces, we may not be able to figure it out any more.
Step2: Trial-and-Error Lattice Attack
@@
We conduct the trial-and-error lattice attack using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz. The parameter of the $block$ size in the BKZ reduction is set to $50$ and the value of $max\_loops$ to $8$. We collect the CPA results of the first sixty $basemuls$ calls. Fig. 4 shows the relationships between the success rate/time and the number of power traces represented in the blue curve and groups of histograms, respectively. The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times. It can be seen that when the number of power traces is no less than $15$, all the experimental attacks are successful. Those cases with traces number smaller than $15$ can be compensated to some degree by running more trials of the lattice attack. Meanwhile, according to the decomposed time histograms in Fig. 4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the traces number is doubled. The main reason could be the different utilization ratios of the multi-threading server. Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack are executed in serial. Therefore, we can summarize that when the success rate is large enough (close/equal to $1$), we can only cost at most $15$ power traces for a guess/call and about $9$ minutes to recover the full key on Kyber512.
Table 1. SCAs on Kyber to Recover a 128-Value Group of $\mathbf{s}$
Work Attack Type No Cip Restriction Attack Target Traces Num for a Call Num of Calls
(xu2021magnifying, ) Pro SPA $\times$ INTT, ME 8$\sim$960 128
(mu2022voltage, ) Pro TA $\times$ MR 11 128
(sim2022chosen, ) Non-Pro SPA $\times$ MR 6$\sim$12 128
(yang2023chosen, ) Non-Pro CPA $\times$ CM 20$\sim$30 128
(kuo2023lattice, ) Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
Ours Non-Pro CPA $\checkmark$ CM 11$\sim$15 60
  • Cip: Ciphertext, Num: Number, Pro: Profile, ME: Message Encoding,
    MR: Modular Reduction, CM: Coefficient Multiplication.
Table 1 shows the comparison results of SCAs on Kyber to recover a 128-value group of $\mathbf{s}$. The attack types include the profiled SPA (xu2021magnifying, ), profiled TA (mu2022voltage, ), non-profiled SPA (sim2022chosen, ), and non-profiled CPA (kuo2023lattice, ; yang2023chosen, ). Our method belongs to the non-profiled CPA, which does not cost extra training step. Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme (ravi2024side, ). The restriction ciphertext CPA requires oracle-based assistance to be applied to the KEM scheme. For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work (yang2023chosen, ) which requires at least $20\times 128=2560$ power traces. If we slightly increase the traces number of the guess, the required time can be reduced to approximate $5$ minutes, much faster than the $20$ minutes recorded in (kuo2023lattice, ).